A 2.03 g sample of a compound consisting of carbon, hydrogen, oxygen, nitrogen, and sulfur was combusted in excess oxygen. This produced 2.19 g CO2 and 1.20 g H2O . A second sample of this compound with a mass of 4.73 g produced 3.10 g SO3 . A third sample of this compound with a mass of 8.37 g produced 4.32 g HNO3 . Determine the empirical formula of the compound. Enter the correct subscripts on the given chemical formula.
empirical formula:
CHNSO
CHNSO
1
Expert's answer
2020-05-27T13:34:13-0400

n(CO2) = 3.63g/44g/mol = 0.0825 mol

n(SO3) = 0.031125 mol

n(HNO3) = 0.046825 mol

M = m/n = 4.89g/0.031125mol = 157 g/mol

Empirical formula of the compound C5H20O2SN